OpenStudy (help!!!!):

13. If two planes leave the same airport at 1:00 PM, how many miles apart will they be at 3:00 PM if one travels directly north at 150 mph and the other travels directly west at 200 mph? A. 50 miles B. 100 miles C. 500 miles D. 700 miles E. 1,000 miles

OpenStudy (anonymous):

Pythgoras Thorem

OpenStudy (anonymous):

500 miles apart. so C

OpenStudy (anonymous):

rignt @Anyssah

OpenStudy (help!!!!):

how did you find that out?

OpenStudy (anonymous):

pythagorean theorem. Since the plane travels north at 150mph for 2 hours, it travelled 300 miles. If the other one traveled west at 200mp for 2 hours, it traveled 400 miles. Pythagorean's theorem states that for right triangles a^2+b^2=c^2 or c=sqrt(a^2+b^2) c=sqrt(400^2+300^2) or 500.

OpenStudy (help!!!!):

wait how did you find the miles?

OpenStudy (anonymous):

Here's how you understand which answer is correct. I find that drawing a picture helps. Both planes will fly for two hours. The first one flies north 300 miles. The second flies east 400 miles. That's a right triangle where the legs are 300 miles and 400 miles. The hypotenuse is how far the planes are apart (shortest distance between their positions). We basically have a 3-4-5 right triangle, but the legs are 300 miles and 400 miles. So the hypotenuse will be 500 miles. But you can use a calculator to check. You would square 300 and add it to the square of 400, then you'd take the square root of that sum. You'll find the answer is 500. So the answer is the planes will be 500 miles apart. Hope that's helpful. It's my first day on the site! :-)

OpenStudy (help!!!!):

thank you everyone, i wish i could give you all a medal